OpenStudy (anonymous):

how to find the modulus and argument of (1-2i)(4+3i).. steps please??

hartnn (hartnn):

first simplify this : (1-2i)(4+3i multiply them out, or foil, and bring it in the form of x+iy can you do that ?

hartnn (hartnn):

once you have it in x+iy form, your modulus will be \(\large \sqrt{x^2+y^2}\) and the argument will be \(arctan (y/x)\)

hartnn (hartnn):

(1-2i)(4+3i) start simplifying :)

hartnn (hartnn):

@Ayuni did you try ? if you're stuck in any step, i could help you.

OpenStudy (marissalovescats):

Welcome to OpenStudy :) @Ayuni

OpenStudy (anonymous):

btw, thanks @marissalovescats

OpenStudy (anonymous):

how about this type of question; \[\frac{ 1+5i }{ 2+i }\] i got the final answer is \[\frac{ 9i+7 }{ i+4 }\]

OpenStudy (anonymous):

@hartnn .. thanks anyway.. helps me a lot.